Becton Labs, Inc., produces various chemical compounds for industrial use. One compound, called Fludex, is prepared using anRequired: 1. For direct materials: a. Compute the price and quantity variances. b. The materials were purchased from a new suReq 1A Req 1B Req 2A Req 2B Reg 3 For direct materials, compute the price and quantity variances. (Indicate the effect favoraReq 1A Req 1B Req 2A Req 2B Reg 3 For direct labor, compute the rate and efficiency variances. (Indicate the effect of e U
Answer #1
Req-1(A)Computation of Material Price & Quanntity Variance
Material price variance = (SP - AP) * AQ purchased
' = ($24 - $22.55) * 11500 = $16675 F
Material quantity variance = (SQ - AQ Used ) * SP
' = (8610 Ounce - 8800 Ounce ) * $24 = $4560 U
Refrence
Standard qty of material = 4100 Units *2.10 Ounces= 8610 ounce
Actual quantity of material purchased = 11500 ounce
Actual quantity of material used = 11500 - 2700 = 8800 ounce
Standard price of material = $24 per ounce
Actual price of material = $259,325 / 11500 = $22.55
Req-1(B),Yes , Company should signed a contract because , Material Price variance shows favourable which shows that actual material procurement price is lesser than Standard Price.
Solution 2a:
Refrence
Standard hours of direct labor = 4100 Unit *0.40 Hour = 1640 hours
Standard rate of direct labor = $14 per hour
Actual hours of direct labor = 23 Technician *100 Hour = 2300 hours
Actual rate of direct labor = $11 per hour
Req-2(a) Labour Rate & Efficiency Variance
Direct labor rate variance = (SR - AR) * AH
= ($14 - $11) * 2300 Hour = $6900 F
Direct labor efficiency variance
= (SH - AH) * SR = (1640 - 2300) * $14= $9240 U
Req:2(b)NO, New labor mix should not be continued because Employing more assistant rather senior technician resulted in favorable direct labor rate variance but unfavorable laor efficiency variance. Further unfavorable efficiency variance is higher than favorable rate variance,
Req-3 Variable OH Rate & Efficiency Variance
Variable overhead rate variance = (SR - AR) * AH
= ($2.50 - $1) *2300 = $3450 F
Variable overhead eficiency variance = (SH - AH) * SR
= (1640 - 2300) * $2.50 = $1650U
Refrence
Standard hours of direct labor = 4100 Unit *0.40 Hour = 1640 hours
Standard rate of variable overhead= $2.50 per hour
Actual hours of direct labor = 23 Technician *100 Hour = 2300 hours
Actual rate of variable overhead = $2300 / 2300 = $1